Search Header Logo
  1. Resource Library
  2. Math
  3. Probability And Statistics
  4. Normal Distribution
  5. Normal Distribution
Normal distribution

Normal distribution

Assessment

Presentation

•

Mathematics

•

11th Grade

•

Practice Problem

•

Medium

•
CCSS
HSS.ID.A.4

Standards-aligned

Created by

Rizki Alfath

Used 14+ times

FREE Resource

17 Slides • 15 Questions

1

Normal Distribution

2

a) Normal Distribution
media

3

a) Normal Distribution
media

4

a) Normal Distribution
media

5

a) Normal Distribution
media

6

a) Normal Distribution
media

7

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the standardized score for a raw score of 70.

1

z = 0

2

z = 1.5

3

z = –0.5

4

z = 2

8

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the standardized score for a raw score of 52.5.

1

z = 0

2

z = 1.5

3

z = –0.5

4

z = 2

9

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the standardized score for a raw score of 66.5.

1

z = 0

2

z = 1.5

3

z = –0.5

4

z = 2

10

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the standardized score for a raw score of 56.

1

z = 0

2

z = 1.5

3

z = –0.5

4

z = 2

11

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the a raw score for standardized score of 1.3.

1

x = 65.1

2

x = 39.2

3

x = 53.2

4

x = 70

12

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the a raw score for standardized score of –2.4.

1

x = 65.1

2

x = 39.2

3

x = 53.2

4

x = 70

13

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the a raw score for standardized score of –0.4.

1

x = 65.1

2

x = 39.2

3

x = 53.2

4

x = 70

14

Multiple Choice

Question image

Given that:

μ = 56 and σ = 7

Find the a raw score for standardized score of 2.0.

1

x = 65.1

2

x = 39.2

3

x = 53.2

4

x = 70

15

b) Standard Normal Distribution
media

16

b) Standard Normal Distribution
media

17

b) Standard Normal Distribution
media

18

b) Standard Normal Distribution
media

19

b) Standard Normal Distribution
media

20

b) Standard Normal Distribution

​If Z ~ N(0, 12) find P(Z < 0.247).

media

​P(Z < 0.247) = 0.5948 + 0.0027 = 0.5975

21

Multiple Choice

If Z ~ N(0, 12) find Φ(1.62).

[Hint: P(Z < 1.62) = Φ(1.62)]

1

0.9474

2

0.2236

3

0.0934

4

0.8438

22

Multiple Choice

If Z ~ N(0, 12) find P(Z > 0.76).

[Hint: P(Z > 0.76) = 1 – Φ(0.76)]

1

0.9474

2

0.2236

3

0.0934

4

0.8438

23

Multiple Choice

If Z ~ N(0, 12) find P(Z < –1.32).

[Hint: P(Z < –1.32) = 1 – Φ(1.32)]

1

0.9474

2

0.2236

3

0.0934

4

0.8438

24

Multiple Choice

If Z ~ N(0, 12) find P(–1.2 < Z < 1.738).

[Hint: P(–1.2 < Z < 1.738) = Φ(1.738) – Φ(–1.2)]

1

0.9474

2

0.2236

3

0.0934

4

0.8438

25

b) Standard Normal Distribution

​If X ~ N(2, 52) find P(X < 7).

media

26

Fill in the Blanks

If X ~ N(2, 52) find P(X < 9).

27

b) Standard Normal Distribution

​If X ~ N(2, 52) find P(X > 11).

media

28

Fill in the Blanks

If X ~ N(2, 52) find P(X > 10).

29

b) Standard Normal Distribution

​If X ~ N(2, 52) find P(|X| < 3).

media

30

Fill in the Blanks

If X ~ N(2, 52) find P(X > 10).

31

b) Standard Normal Distribution

​If X ~ N(2, 52) find P(|X - 2| < 6).

media

32

b) Standard Normal Distribution

​If X ~ N(2, 52).

Find x such that P(X > x) = 0.05

media
media
pattern-tertiary
Normal Distribution

Show answer

Auto Play

Slide 1 / 32

SLIDE